OpenStudy (anonymous):

Determine the domain of the function h(x) = 8x / x(x^2-81)

OpenStudy (nurali):

h(x) = 8x/ x(x^2-81). That is h(x) = 8/(x^2-81) (canceling x not zero) Since division by zero is not defined, h(x) cannot be defined for (x^2-81)=0 And (x^2-81) = 0 implies (x+9)(x-9) = 0 (using a^2-b^2 = (a+b)(a-b) where a = x and b = 9) that is for (x+9)=0 giving x=-9 and for (x-9) = 0 giving x = 9 Therefore the funciton h(x) is defined for all x EXCEPTING x = -9 and x = 9 Therefore the funciton h(x) is defined for all x other than -9 and 9 Therefore the domain of the given function is the complement of {-9,9} in the real number system R That is Domain = R - {-9,9}

OpenStudy (anonymous):

thank you!

OpenStudy (nurali):

My Pleasure.